Throwing a dice 3 times, where X the combination of total values of 13. What is all possible combinations of value that X can take
Answer #1

A dice is rolled 3 times. A roll can take values 1,2, 3, 4, 5,6. We need the sum of 3 rolls to be 13. We enumerate the outcomes that give a sum of 13 for the 3 dies. Let the outcome of the rolls be 11-12 , ЛА .

Then with X_1=1 requires X_2+X_3=12Rightarrow X_2=6,X_3=6 , 1 combination

Then with X_1=2 requires X_2+X_3=11Rightarrow X_2=6,X_3=5 ;X_2=5,X_3=6 2 combinations

Then with X_1=3 requires X_2+X_3=10Rightarrow X_2=6,X_3=4 ;X_2=5,X_3=5;X_2=4,X_3=6 3 combinations

Then with X_1=4 requires X_2+X_3=9Rightarrow X_2=6,X_3=3 ;X_2=5,X_3=4;X_2=4,X_3=5;X_2=4,X_3=5;X_2=3,X_3=6 8 combinations

Then with X_1=5 requires X_2+X_3=8Rightarrow X_2=6,X_3=2;X_2=5,X_3=3;X_2=4,X_3=4;X_2=5,X_3=3;X_2=2,X_3=6 5 combinations

Then with X_1=6 requires X_2+X_3=7Rightarrow X_2=6,X_3=1; X_2=5,X_3=2; X_2=4,X_3=3; X_2=3,X_4=3; X_2=2,X_3=5; X_2=1,X_3=6 6 combinations

Thus the total combinations of values X can take is 1+2+3+8+6={color{Blue} 22} .